Construct a combinational circuit using 2-input basic gates which has 3 inputs A,B & C

and 3 outputs X,Y &Z. When the decimal equivalent of the binary input is 0, 1, 5 or 3,

the decimal equivalent of the binary output is one greater than the input and when

decimal equivalent of the binary input is 4,2,6 &7 , the decimal equivalent of binary

output is one less than the input.

1
Expert's answer
2021-09-26T11:56:14-0400

Design procedure:

1. Derive the truth table that defines the required relationship between

inputs and outputs.


Obtain the simplified Boolean functions for each output as a function of the

input variables.

Map for output A:

The simplified expression from the map is:

"A =xz+xy+yz"



Map for output B:

The simplified expression from the map is:

"B = x'y'z +x'yz'+ xy'z'+ xyz"



Map for output C:

The simplified expression from the map is:

C =z'



Draw the logic diagram.

"A =xz+xy+yz\\
B = x'y'z +x'yz'+ xy'z'+ xyz\\
C=z'"